# FedQUIT: On-Device Federated Unlearning via a Quasi-Competent Virtual Teacher

### v2

@@ 5.2 Metrics
Accuracy on forgetting data. We monitor the unlearned model’s accuracy on client $u$’s train data, and we track the gap with the retrained model [21]. This metric provides insight into how closely the unlearning process approximates the desired outcome achieved through retraining.
Membership Inference Attacks (MIA). To assess unlearning, we also use the MIA success rate. MIAs attempt to determine whether an individual sample was included in the training set of a model. The success rate of these attacks indicates how many data samples in $D_{u}$ can be correctly predicted as training records of $w_{u}$. We consider as most effective the approach that minimizes the absolute difference with the retrained model’s MIA value. We consider two different MIAs: (1) a confidence-based MIA predictor [29] and (2) a loss-based MIA predictor [36]. Appendix 12 provides detailed description of the attacks.
Recovery rounds. FU mechanisms need a recovery period (recovery phase) to match the performance of the retrained model [27]. For the federated settings, we monitor how many communication rounds the unlearned model needs to match or surpass the accuracy on the test data (referred to as Test Acc.) of the retrained model. Note that the reported Test Acc. for FU baselines is higher by design, as recovery ends once it exceeds the retrained model’s accuracy.
@@ 5.3 Main Results
Algorithm Rounds ($\downarrow$) CE ($\uparrow$) Test Acc. Forget Acc. Original 75.03 ±0.0 84.25 ±5.6 Retrained 73.30 ±0.8 57.80 ±5.3 PGA [12] 6.9 ±4.3 7.2$\times$ 73.60 ±0.4 62.15 (4.35 ±3.3) FedQUIT 8.7 ±2.5 5.7$\times$ 73.42 ±0.9 58.50 (1.86 ±1.5)
@@
First, we evaluate FedQUIT in federated settings, comparing it with two baselines: the Incompetent Teacher baseline (Section 2) and the state-of-the-art PGA baseline [12]. Table 3 presents results on federated CIFAR-100 and CIFAR-10 using ResNet-18 as the model architecture, across different levels of data heterogeneity and local epochs. Table 4 reports results for federated CIFAR-100 with a visual transformer (extended results in Appendix Table 9). In Appendix Table 11, we also report results for the Natural baseline [27] as a reference, which corresponds to simply excluding the client that should be forgotten from the federation without applying any unlearning method, demonstrating that active unlearning is needed.
@@
While FedQUIT’s communication cost for the unlearning round is the same as FedAvg, FedQUIT significantly reduces cumulative communication costs over retraining, as shown in the CE column of Table 3. In fact, it achieves 20$\times$ reduction with IID data and up to 117.6$\times$ reduction with non-IID data. FedQUIT also outperforms PGA in this regard, with the only exception being the case MiT-B0/CIFAR-100 (see Table 4 and Appendix Table 9). In this setting, PGA shows a slightly higher gain but fails to match the retrained model’s forgetting metrics, with a higher delta in Forget Acc. indicating a less effective forgetting.
@@ 6.2 Forget Robustness over Fine Tuning
For completeness, we evaluate the robustness of FedQUIT over additional epochs of fine tuning in a centralized setting, i.e., whether the forgetting metrics remain stable over time after unlearning or if the influence of the removed data reemerges. We use a straightforward centralized version of FedQUIT (see Sec. 13 for detail), and we compare it against the state-of-the-art MU baseline from [5] (see extended discussion in Sec. 13). We replicate the CIFAR-Super20 setup from [5], which restructures CIFAR-100’s 60,000 images into 20 super-classes, each containing five original CIFAR-100 classes (e.g., the people super-class includes baby, boy, girl, man, woman). In the experiments, we unlearn all examples of a specific sub-class (e.g., baby), making unlearning challenging since removing one sub-class may affect other visually similar samples within the same super-class. We evaluate five cases, unlearning the sub-classes baby, rocket, lamp, sea, and mushroom. Alongside the key metrics from [5], we track test accuracy on adjacent data—samples from the same super-class as the unlearned sub-class but excluding the unlearned sub-class itself. This helps assess whether the methods can selectively erase the contribution of the forgetting data without degrading related information. ResNet-18 is used as the neural network, with images resized to 224×224 and normalized.
Table 6 presents results after one epoch of fine tuning on retain data. Appendix Tables 13 - 17 provide extended results also after 2 and 3 fine-tuning epochs. While the method from [5] produces an unlearned model with retain data accuracy comparable to retrained models even before fine-tuning, it undergoes significant degradation in forget metrics when fine-tuned further on retain data. This is evident from the high positive accuracy gap on forget data after just one fine-tuning epoch, such as +34% in the lamp case. In contrast, FedQUIT preserves forget metrics near the gold standard, with only a slight increase over the second and third fine-tuning epochs, demonstrating its effectiveness and robustness in forgetting.
Task Algorithm Retain Data Adjacent Data Forget Data MIA [29] Baby Original 85.68 94.92 91.00 98.80 Retrain 85.92 ±0.21 93.64 ±0.12 82 ±0.41 73.6 ±0.22 FedQUIT (central.) 85.01 ±0.23(-0.92) 88.83 ±4.07(-4.81) 80.33 ±4.04(-1.67) 67 ±6.22(-6.6) Chundawat et al. 84.40 ±0.3(-1.53) 91.24 ±0.62(-2.4) 82 ±2(0) 77.67 ±3.21(4.07) Rocket Original 85.57 91.04 92.00 99.40 Retrain 85.39 ±0.26 81.4 ±0.32 3 ±0.34 0.8 ±0.29 FedQUIT (central.) 85.16 ±0.08(-0.23) 85.22 ±4.72(3.82) 7 ±2(4) 1.07 ±0.58(0.27) Chundawat et al. 84.90 ±0.11(-0.48) 81.94 ±1.77(0.54) 25.33 ±7.02(22.33) 6.47 ±3.8(5.67) Lamp Original 85.03 86.11 78.00 88.60 Retrain 84.93 ±0.46 90.76 ±0.31 15 ±0.18 8.2 ±0.35 FedQUIT (central.) 85.07 ±0.32(0.14) 87.64 ±4.21(-3.12) 18 ±10.54(3) 8.13 ±4.39(-0.07) Chundawat et al. 84.62 ±0.32(-0.32) 89.52 ±3.06(-1.24) 49 ±3(34) 26.93 ±4.41(18.73)
@@ 8 Inspiring Observations
In FL, and especially in cross-device FL, data are supposed to be non-IID. This characteristic translates to drift of client models, that diverge from the global model making the next FedAvg aggregation inefficient. Extensive research has investigated the use of the global model as an anchor for local models to mitigate this drift, aiming to minimize the disruptive impact of data heterogeneity on aggregation. An approach to implementing this guidance is via KD, with the global model acting as the teacher and the local model as the student [35, 19, 23, 10]. This approach preserves global knowledge while incorporating local adaptations, which helps limit the phenomenon of catastrophic forgetting – the rapid overwriting of previous knowledge to accommodate the distribution shift imposed by the specific clients participating in each round.
@@ 9 Similarly Inspired Work in FL
The mechanisms that we present in this paper use a student-teacher framework locally at FL clients to retain the good knowledge from the original model while selectively scrubbing the contributions to forget. We use a crafted version of the FL global model as the teacher, serving as a natural proxy for the retain data that cannot be directly accessed in FL. Recently, KD [14, 3] has been widely used in FL to address various issues [25], such as reducing communication costs, enabling model heterogeneity within the federation, improving generalization or personalization performance with heterogeneous data, and even performing unlearning. Most relevantly for this paper, numerous works employ the global model as a KD teacher for the local model to mitigate catastrophic forgetting and client drift in the presence of heterogeneous data (e.g., [35, 19, 23, 10]). For example, FedGKD [35] uses the global model output as a guide for local models, limiting client model drifts. FedNTD [19] demonstrated that excluding the teacher’s true-class logit during the distillation phase is beneficial for local-global distillation. Note that FedNTD [19] and FedQUIT differ significantly in both objective and methodology: FedNTD addresses data heterogeneity, while FedQUIT enables client unlearning. Technically, (1) FedNTD applies KL-divergence as a local regularizer for cross-entropy loss, whereas FedQUIT directly optimizes the KL divergence in Eq. 2; (2) FedNTD ignores the logits produced by the true classes, while FedQUIT actively penalizes them—an essential step to induce forgetting. Other recent contributions have proposed other label-masking approaches for the global model’s output (i.e., the teacher) [23, 10]. Unlike these works, our approach modifies the output of the global model, which still acts as a teacher for the local model, to perform unlearning.
@@ 10 Extended Description of Table 1
Historical information. This refers to whether a method requires storing and accessing historical data, such as the complete history of per-client updates, which is typically maintained by the parameter server. Note that: (1) Linking per-client update histories to specific clients requesting unlearning undermines FL’s privacy design; (2) to some extent it violates the ephemeral updates requirement of the FL framework [16]; and (3) at scale, considering a massive number of clients [6], several FL rounds, and multiple simultaneous learning tasks, the corresponding storage requirements might become impractical.
@@ 11 Infrastructure and Libraries
Figure 6: Label distribution across clients (0-9) for CIFAR-10 (Non-IID, $\alpha=0.3$) and CIFAR-100 (IID and Non-IID,$\alpha=0.1$).
@@ 12 Membership Inference Attacks
Song’s MIA [29].
This attack is implemented using the prediction confidence-based attack proposed in [29]. This approach includes a training phase where a balanced dataset is sampled from data seen during training and data unseen during traning. For seen data, we use the retain dataset, which includes the training data but excludes any forget data; in federated settings, this retain dataset consists of the combined data from the remaining clients. For unseen data, we use a standard test dataset that the model has never encountered during training and is distinct from the forgetting dataset. The MIA predictor is trained to identify whether the target model’s output resembles that of known (seen) data or unknown (unseen) data. To assess unlearning effectiveness, the MIA predictor is then applied to a model using the forgetting data as input.
Yeom’s MIA [36].
@@ 13 Baselines in Experiments
PGA [12] (federated baseline):
The work in [12] applies projected gradient ascent (PGA) to eliminate the influence of the data to be forgotten, directly on-device (similarly to our method). PGA constrains the local weight updates to ensure that the model weights remain within an $l2$-norm ball around a carefully selected reference model. The reference model is computed by removing the last unlearning client’s model update from the current global model. Halimi et al. perform early stopping if the distance between the unlearned model and the reference model is below a threshold, which must be tuned. Note that: (1) PGA requires clients to store their last model updates to build the so-called reference model, and, to do that, it maintains a state on clients; (2) in light of the previous point, PGA needs full participation of clients; (3) PGA needs to tune two thresholds (plus the standard training hyper-parameters).
Competent-Incompetent Teacher [5] (centralized baseline):
During the unlearning process, forget data and retain data are interleaved. The model is trained to mimic the output of the original model when presented with retain data, while it emulates the output of an incompetent model (e.g., a random model) when given forget data. It is worth noting that forget data and retain data need to be properly balanced through an additional hyper-parameter that controls the proportion of the two sets. Note that FedQUIT is foundamentally different from this approach: (1) FedQUIT explicitly targets FL settings and not centralized settings, even though it can be applied in the latter setting as well (as we demonstrated in the Experimental Results Section), (2) FedQUIT does not need retain data (which are not available in FL settings when working directly on the unlearning client’s device), (3) FedQUIT tailors a virtual teacher and modifies the output of the FL global model to induce unlearning, (4) The work in [5] also needs to tune a ratio to balance the amount retain data and forget data.
@@ 14.1 Regular Training (Federated Settings)
We used a standard ResNet-18 [13] and employed Group Normalization layers, similar to other works with similar settings (e.g., [17]). We used SGD as a local optimizer with a learning rate set to 0.1, with a round-wise exponential decay of 0.998, 1 local epoch, local batch size of 32. We pre-processed the training images with random crop, horizontal flip and normalization layers. During unlearning routines, we only apply normalization. As server-side optimizer, we consider SGD with a learning rate set to 1.0 (standard FedAvg aggregation and update application).
@@
We used a visual transformer, i.e., MiT-B0 [33], with approximately 3.6M parameters, initialized from a pre-trained model checkpoint trained on ImageNet-1k (69.27% accuracy on test data). We adapted the one-layer classification head to this task, initializing such a layer from scratch. We employed the AdamW optimizer with a client learning rate of 3e-4, with a round-wise exponential decay of 0.998, 1 local epoch, local batch size of 32, and weight decay regularization of 1e-3. As server-side optimizer, we consider SGD with a learning rate set to 1.0 (standard FedAvg aggregation and update application). The images are resized to a resolution of 224x224.
@@ 14.2 Unlearning Phase (Federated Settings)
PGA [12]: the same hyper-parameter prescribed in the original paper, and we tuned their early stopping threshold in $[2.2,10.0]$ selecting the following as best: 9.0 (IID), 6.5 (Non-IID).
@@
FedQUIT (all variants), Incompetent Teacher: 1 unlearning epochs, learning rate 5e-4, AdamW optimizer, local batch size 32. PGA [12]: we report in Table 8 the best configs of hyperparameter (learning rate for local unlearning $\mu_{PGA}$, optimizer opt, unlearning epochs Eu, threshold $\tau$). We selected ($\mu_{PGA}$=3e-4, opt=AdamW, Eu=5, $\tau=7.5$) as best. We set batch size to 512 (similar to what prescribed in the original paper).
Algorithm Rounds $\Delta$ Forget Acc. $\Delta$ Test Acc. $\Delta$ MIA [29] PGA (1e-2, SGDM, 5, 6.5) 6.4±7.5 10.52±7.88 0.62±0.77 11.54±9.25 PGA (1e-2, SGDM 5, 7.0) 12.3±11.09 8.98±7.19 0.56±0.8 9.62±9.01 PGA (3e-4, AdamW, 5, 7.0) 3.8±2.97 8.61±6.64 0.65±0.85 8.24±7.23 PGA (3e-4, AdamW, 5, 7.5) 6.9±4.28 4.35±3.28 0.25±0.36 4.98±3.67 PGA (3e-4, AdamW, 5, 8.0) 7.5±3.74 5.06±3.11 0.33±0.27 5.81±3.67
@@
For the baseline from [5] we used the hyperparameters from the original paper. For FedQUIT (all variants) we used the Adam optimizer with learning rate of 5e-4 (tuned in $[1e-1,1e-2,1e-3,5e-4,1e-4,1e-5]$).

### v3

@@ 5.2 Metrics
Accuracy on forgetting data. We monitor the unlearned model’s accuracy on client $u$’s train data, and we track the gap with the retrained model [liu2024model]. This metric provides insight into how closely the unlearning process approximates the desired outcome achieved through retraining.
Membership Inference Attacks (MIA). To assess unlearning, we also use the MIA success rate. MIAs attempt to determine whether an individual sample was included in the training set of a model. The success rate of these attacks indicates how many data samples in $D_{u}$ can be correctly predicted as training records of $w_{u}$. We consider as most effective the approach that minimizes the absolute difference with the retrained model’s MIA value. We consider two different MIAs: (1) a confidence-based MIA predictor [song2021systematic] and (2) a loss-based MIA predictor [yeom2018privacy]. Appendix 12 provides detailed description of the attacks.
Recovery rounds. FU mechanisms need a recovery period (recovery phase) to match the performance of the retrained model [romandini2024federated]. For the federated settings, we monitor how many communication rounds the unlearned model needs to match or surpass the accuracy on the test data (referred to as Test Acc.) of the retrained model. Note that the reported Test Acc. for FU baselines is higher by design, as recovery ends once it exceeds the retrained model’s accuracy.
@@ 5.3 Main Results
[…]
@@
First, we evaluate FedQUIT in federated settings, comparing it with two baselines: the Incompetent Teacher baseline (Section 2) and the state-of-the-art PGA baseline [halimi2022federated]. Table 3 presents results on federated CIFAR-100 and CIFAR-10 using ResNet-18 as the model architecture, across different levels of data heterogeneity and local epochs. Table 4 reports results for federated CIFAR-100 with a visual transformer (extended results in Appendix Table 9). In Appendix Table 11, we also report results for the Natural baseline [romandini2024federated] as a reference, which corresponds to simply excluding the client that should be forgotten from the federation without applying any unlearning method, demonstrating that active unlearning is needed.
@@
[…]
@@ 6.2 Forget Robustness over Fine Tuning
For completeness, we evaluate the robustness of FedQUIT over additional epochs of fine tuning in a centralized setting, i.e., whether the forgetting metrics remain stable over time after unlearning or if the influence of the removed data reemerges. We use a straightforward centralized version of FedQUIT (see Sec. 13 for detail), and we compare it against the state-of-the-art MU baseline from [chundawat2023can] (see extended discussion in Sec. 13). We replicate the CIFAR-Super20 setup from [chundawat2023can], which restructures CIFAR-100’s 60,000 images into 20 super-classes, each containing five original CIFAR-100 classes (e.g., the people super-class includes baby, boy, girl, man, woman). In the experiments, we unlearn all examples of a specific sub-class (e.g., baby), making unlearning challenging since removing one sub-class may affect other visually similar samples within the same super-class. We evaluate five cases, unlearning the sub-classes baby, rocket, lamp, sea, and mushroom. Alongside the key metrics from [chundawat2023can], we track test accuracy on adjacent data—samples from the same super-class as the unlearned sub-class but excluding the unlearned sub-class itself. This helps assess whether the methods can selectively erase the contribution of the forgetting data without degrading related information. ResNet-18 is used as the neural network, with images resized to 224×224 and normalized.
Table 6 presents results after one epoch of fine tuning on retain data. Appendix Tables 13 - 17 provide extended results also after 2 and 3 fine-tuning epochs. While the method from [chundawat2023can] produces an unlearned model with retain data accuracy comparable to retrained models even before fine-tuning, it undergoes significant degradation in forget metrics when fine-tuned further on retain data. This is evident from the high positive accuracy gap on forget data after just one fine-tuning epoch, such as +34% in the lamp case. In contrast, FedQUIT preserves forget metrics near the gold standard, with only a slight increase over the second and third fine-tuning epochs, demonstrating its effectiveness and robustness in forgetting.
Task Algorithm Retain Data Adjacent Data Forget Data MIA [song2021systematic] Baby Original 85.68 94.92 91.00 98.80 Retrain 85.92 ±0.21 93.64 ±0.12 82 ±0.41 73.6 ±0.22 FedQUIT (central.) 85.01 ±0.23(-0.92) 88.83 ±4.07(-4.81) 80.33 ±4.04(-1.67) 67 ±6.22(-6.6) chundawat2023can 84.40 ±0.3(-1.53) 91.24 ±0.62(-2.4) 82 ±2(0) 77.67 ±3.21(4.07) Rocket Original 85.57 91.04 92.00 99.40 Retrain 85.39 ±0.26 81.4 ±0.32 3 ±0.34 0.8 ±0.29 FedQUIT (central.) 85.16 ±0.08(-0.23) 85.22 ±4.72(3.82) 7 ±2(4) 1.07 ±0.58(0.27) chundawat2023can 84.90 ±0.11(-0.48) 81.94 ±1.77(0.54) 25.33 ±7.02(22.33) 6.47 ±3.8(5.67) Lamp Original 85.03 86.11 78.00 88.60 Retrain 84.93 ±0.46 90.76 ±0.31 15 ±0.18 8.2 ±0.35 FedQUIT (central.) 85.07 ±0.32(0.14) 87.64 ±4.21(-3.12) 18 ±10.54(3) 8.13 ±4.39(-0.07) chundawat2023can 84.62 ±0.32(-0.32) 89.52 ±3.06(-1.24) 49 ±3(34) 26.93 ±4.41(18.73)
@@ 8 Inspiring Observations
In FL, and especially in cross-device FL, data are supposed to be non-IID. This characteristic translates to drift of client models, that diverge from the global model making the next FedAvg aggregation inefficient. Extensive research has investigated the use of the global model as an anchor for local models to mitigate this drift, aiming to minimize the disruptive impact of data heterogeneity on aggregation. An approach to implementing this guidance is via KD, with the global model acting as the teacher and the local model as the student [yao2021local, lee2021preservation, lu2023federated, guo2024not]. This approach preserves global knowledge while incorporating local adaptations, which helps limit the phenomenon of catastrophic forgetting – the rapid overwriting of previous knowledge to accommodate the distribution shift imposed by the specific clients participating in each round.
@@ 9 Similarly Inspired Work in FL
The mechanisms that we present in this paper use a student-teacher framework locally at FL clients to retain the good knowledge from the original model while selectively scrubbing the contributions to forget. We use a crafted version of the FL global model as the teacher, serving as a natural proxy for the retain data that cannot be directly accessed in FL. Recently, KD [hinton2015distilling, bucilua2006model] has been widely used in FL to address various issues [mora2022knowledge], such as reducing communication costs, enabling model heterogeneity within the federation, improving generalization or personalization performance with heterogeneous data, and even performing unlearning. Most relevantly for this paper, numerous works employ the global model as a KD teacher for the local model to mitigate catastrophic forgetting and client drift in the presence of heterogeneous data (e.g., [yao2021local, lee2021preservation, lu2023federated, guo2024not]). For example, FedGKD [yao2021local] uses the global model output as a guide for local models, limiting client model drifts. FedNTD [lee2021preservation] demonstrated that excluding the teacher’s true-class logit during the distillation phase is beneficial for local-global distillation. Note that FedNTD [lee2021preservation] and FedQUIT differ significantly in both objective and methodology: FedNTD addresses data heterogeneity, while FedQUIT enables client unlearning. Technically, (1) FedNTD applies KL-divergence as a local regularizer for cross-entropy loss, whereas FedQUIT directly optimizes the KL divergence in Eq. 2; (2) FedNTD ignores the logits produced by the true classes, while FedQUIT actively penalizes them—an essential step to induce forgetting. Other recent contributions have proposed other label-masking approaches for the global model’s output (i.e., the teacher) [lu2023federated, guo2024not]. Unlike these works, our approach modifies the output of the global model, which still acts as a teacher for the local model, to perform unlearning.
@@ 10 Extended Description of Table 1
Historical information. This refers to whether a method requires storing and accessing historical data, such as the complete history of per-client updates, which is typically maintained by the parameter server. Note that: (1) Linking per-client update histories to specific clients requesting unlearning undermines FL’s privacy design; (2) to some extent it violates the ephemeral updates requirement of the FL framework [kairouz2019advances]; and (3) at scale, considering a massive number of clients [dean2012large], several FL rounds, and multiple simultaneous learning tasks, the corresponding storage requirements might become impractical.
@@ 11 Infrastructure and Libraries
[…]
@@ 12 Membership Inference Attacks
Song’s MIA [song2021systematic].
This attack is implemented using the prediction confidence-based attack proposed in [song2021systematic]. This approach includes a training phase where a balanced dataset is sampled from data seen during training and data unseen during traning. For seen data, we use the retain dataset, which includes the training data but excludes any forget data; in federated settings, this retain dataset consists of the combined data from the remaining clients. For unseen data, we use a standard test dataset that the model has never encountered during training and is distinct from the forgetting dataset. The MIA predictor is trained to identify whether the target model’s output resembles that of known (seen) data or unknown (unseen) data. To assess unlearning effectiveness, the MIA predictor is then applied to a model using the forgetting data as input.
Yeom’s MIA [yeom2018privacy].
@@ 13 Baselines in Experiments
PGA [halimi2022federated] (federated baseline):
The work in [halimi2022federated] applies projected gradient ascent (PGA) to eliminate the influence of the data to be forgotten, directly on-device (similarly to our method). PGA constrains the local weight updates to ensure that the model weights remain within an $l2$-norm ball around a carefully selected reference model. The reference model is computed by removing the last unlearning client’s model update from the current global model. halimi2022federated perform early stopping if the distance between the unlearned model and the reference model is below a threshold, which must be tuned. Note that: (1) PGA requires clients to store their last model updates to build the so-called reference model, and, to do that, it maintains a state on clients; (2) in light of the previous point, PGA needs full participation of clients; (3) PGA needs to tune two thresholds (plus the standard training hyper-parameters).
Competent-Incompetent Teacher [chundawat2023can] (centralized baseline):
During the unlearning process, forget data and retain data are interleaved. The model is trained to mimic the output of the original model when presented with retain data, while it emulates the output of an incompetent model (e.g., a random model) when given forget data. It is worth noting that forget data and retain data need to be properly balanced through an additional hyper-parameter that controls the proportion of the two sets. Note that FedQUIT is foundamentally different from this approach: (1) FedQUIT explicitly targets FL settings and not centralized settings, even though it can be applied in the latter setting as well (as we demonstrated in the Experimental Results Section), (2) FedQUIT does not need retain data (which are not available in FL settings when working directly on the unlearning client’s device), (3) FedQUIT tailors a virtual teacher and modifies the output of the FL global model to induce unlearning, (4) The work in [chundawat2023can] also needs to tune a ratio to balance the amount retain data and forget data.
@@ 14.1 Regular Training (Federated Settings)
We used a standard ResNet-18 [he2016deep] and employed Group Normalization layers, similar to other works with similar settings (e.g., [kim2022multi]). We used SGD as a local optimizer with a learning rate set to 0.1, with a round-wise exponential decay of 0.998, 1 local epoch, local batch size of 32. We pre-processed the training images with random crop, horizontal flip and normalization layers. During unlearning routines, we only apply normalization. As server-side optimizer, we consider SGD with a learning rate set to 1.0 (standard FedAvg aggregation and update application).
@@
We used a visual transformer, i.e., MiT-B0 [xie2021segformer], with approximately 3.6M parameters, initialized from a pre-trained model checkpoint trained on ImageNet-1k (69.27% accuracy on test data). We adapted the one-layer classification head to this task, initializing such a layer from scratch. We employed the AdamW optimizer with a client learning rate of 3e-4, with a round-wise exponential decay of 0.998, 1 local epoch, local batch size of 32, and weight decay regularization of 1e-3. As server-side optimizer, we consider SGD with a learning rate set to 1.0 (standard FedAvg aggregation and update application). The images are resized to a resolution of 224x224.
@@ 14.2 Unlearning Phase (Federated Settings)
PGA [halimi2022federated]: the same hyper-parameter prescribed in the original paper, and we tuned their early stopping threshold in $[2.2,10.0]$ selecting the following as best: 9.0 (IID), 6.5 (Non-IID).
@@
FedQUIT (all variants), Incompetent Teacher: 1 unlearning epochs, learning rate 5e-4, AdamW optimizer, local batch size 32. PGA [halimi2022federated]: we report in Table 8 the best configs of hyperparameter (learning rate for local unlearning $\mu_{PGA}$, optimizer opt, unlearning epochs Eu, threshold $\tau$). We selected ($\mu_{PGA}$=3e-4, opt=AdamW, Eu=5, $\tau=7.5$) as best. We set batch size to 512 (similar to what prescribed in the original paper).
Algorithm Rounds $\Delta$ Forget Acc. $\Delta$ Test Acc. $\Delta$ MIA [song2021systematic] PGA (1e-2, SGDM, 5, 6.5) 6.4±7.5 10.52±7.88 0.62±0.77 11.54±9.25 PGA (1e-2, SGDM 5, 7.0) 12.3±11.09 8.98±7.19 0.56±0.8 9.62±9.01 PGA (3e-4, AdamW, 5, 7.0) 3.8±2.97 8.61±6.64 0.65±0.85 8.24±7.23 PGA (3e-4, AdamW, 5, 7.5) 6.9±4.28 4.35±3.28 0.25±0.36 4.98±3.67 PGA (3e-4, AdamW, 5, 8.0) 7.5±3.74 5.06±3.11 0.33±0.27 5.81±3.67
@@
For the baseline from [chundawat2023can] we used the hyperparameters from the original paper. For FedQUIT (all variants) we used the Adam optimizer with learning rate of 5e-4 (tuned in [1e1,1e2,1e3,5e4,1e4,1e5][1e-1,1e-2,1e-3,5e-4,1e-4,1e-5]).